# Model Predictive Inferential Control of Neural State-Space Models for Autonomous Vehicle Motion Planning

### v5

@@ VIII Motion Planning for a Real-World Vehicle
The considered vehicle is a full-size Roborace DevBot 2.0 racing car, developed at the Technical University of Munich [63, 75]. The car is equipped with a suite of sensors to measure the longitudinal/lateral velocity/acceleration, yaw rate, steering angle, and other parameters. Data were collected from the car driving on different race tracks at high speeds and under varying tire-road friction scenarios. As shown in [63], end-to-end neural networks are more accurate in capturing the car’s dynamic behavior than the single-track model, especially in highly dynamic situations, given the availability of sufficient data. Also, neural network-based modeling can easily accommodate different tire-road friction levels, while physics-based modeling either fails or requires manual calibration. For this car, high-fidelity dynamics simulation data are publicly available at [76] and used here to train an NSS model for motion planning here.

### v6

@@ VIII Motion Planning for a Real-World Vehicle
The considered vehicle is a full-size Roborace DevBot 2.0 racing car, developed at the Technical University of Munich [Herman:CIST:2020, IAC:JFR:2023]. The car is equipped with a suite of sensors to measure the longitudinal/lateral velocity/acceleration, yaw rate, steering angle, and other parameters. Data were collected from the car driving on different race tracks at high speeds and under varying tire-road friction scenarios. As shown in [Herman:CIST:2020], end-to-end neural networks are more accurate in capturing the car’s dynamic behavior than the single-track model, especially in highly dynamic situations, given the availability of sufficient data. Also, neural network-based modeling can easily accommodate different tire-road friction levels, while physics-based modeling either fails or requires manual calibration. For this car, high-fidelity dynamics simulation data are publicly available at [TUM:GitHub:2021] and used here to train an NSS model for motion planning here.
@@ IX Conclusion
[Uncaptioned image] Iman Askari (Graduate Student Member, IEEE) received the B.S. degree in 2018 and is currently working towards a Ph.D. degree, both in Mechanical Engineering and at the University of Kansas, Lawrence, KS, USA. His research interests lie at the intersections of control theory, estimation theory, and machine learning for robotics applications.
[Uncaptioned image] Ali Vaziri (Graduate Student Member, IEEE) received the B.S. degree in Marine Engineering from Sharif University of Technology, Tehran, Iran, in 2021, where he conducted research on motion planning for surface marine robots. He is currently pursuing a Ph.D. degree in Mechanical Engineering at the University of Kansas, Lawrence, KS, USA. His current research interests include estimation theory and optimal control of high-dimensional data-driven dynamical models.
[Uncaptioned image] Xumein Tu received the B.S. degree in Computer Sciences from Beijing Normal University, China in 1997, the M.S. degree in Applied Mathematics from Worcester Polytechnic Institute in 2002, and the Ph.D. degree in Mathematics in 2006 from New York University. She is currently a Professor at the Department of Mathematics, University of Kansas, Lawrence, KS, USA. Her main research interests are numerical partial differential equations and data assimilations.
[Uncaptioned image] Shen Zeng (Member, IEEE) studied Engineering Cybernetics, Mechatronics, and Mathematics at the University of Stuttgart, from where he also received a Ph.D. degree in 2016. He is currently an Assistant Professor at Washington University in St. Louis in the Department of Electrical and Systems Engineering. His research interests are in systems theory, and, more generally, applied mathematics.
[Uncaptioned image] Huazhen Fang (Member, IEEE) received the B.Eng. degree in Computer Science and Technology from Northwestern Polytechnic University, Xi’an, China, in 2006, the M.Sc. degree from the University of Saskatchewan, Saskatoon, SK, Canada, in 2009, and the Ph.D. degree from the University of California, San Diego, CA, USA, in 2014, both in mechanical engineering. He is currently an Associate Professor of Mechanical Engineering with the University of Kansas, Lawrence, KS, USA. His research interests include control and estimation theory with application to energy management and robotics. He was a recipient of the 2019 National Science Foundation CAREER Award. He currently serves as an Associate Editor for IEEE Transactions on Industrial Electronics and IEEE Control Systems Letters.